Analysis of null.tex.aux, according to rules of Physical Review: All bibitems must occur in the bibliography/reference section in order of their first citation in the file, and all bibitems must be cited in the paper.

* Bibitems that were not cited: none.

* Undefined citations: none.

* The order of the bibitems is OK.